OpenStudy (anonymous):

The rational function \(c(x) := \frac{x+1}{x-1}\) is called Cayley transformation b) Show that \(c:U\rightarrow V\) is bijective

OpenStudy (blockcolder):

Are \(U, V \subseteq \mathbb{R}\)?

OpenStudy (anonymous):

hmm thats not given but above its written there rational function c(x):... i think its U, V \subset \mathbb{R}

OpenStudy (anonymous):

also i mean we suppose yes..

OpenStudy (anonymous):

understand?

OpenStudy (blockcolder):

First show that \(c(x)\) is one-to-one. \(c(x)\) is one-to-one if for all \(u_1, u_2\in U, c(u_1)=c(u_2) \Rightarrow u_1=u_2\). Then show that \(c(x)\) is onto. \(c(x)\) is onto if for all \(v\in V, \exists u\in U(c(u)=v)\).

OpenStudy (anonymous):

ok thank you blockfolder i have 2 more smilar question, i will post it very soon. i would be happy if you take look at it too